OpenStudy (anonymous):

find the vertex and zeros on the parabola y=2x^2+3x-2?

OpenStudy (anonymous):

do i use the formula x=-b/2a for the vertex?

OpenStudy (anonymous):

I need an explanation, not the answer.

OpenStudy (anonymous):

The vertex is: \[x = (\frac{ -3 }{ 4 } , \frac{ -25 }{ 8 })\]